# Relightable Neural Actor with Intrinsic Decomposition and Pose Control

### v1

@@ Abstract
Creating a digital human avatar that is relightable, drivable, and photorealistic is a challenging and important problem in Vision and Graphics. Humans are highly articulated creating pose-dependent appearance effects like self-shadows and wrinkles, and skin as well as clothing require complex and space-varying BRDF models. While recent human relighting approaches can recover plausible material-light decompositions from multi-view video, they do not generalize to novel poses and still suffer from visual artifacts. To address this, we propose Relightable Neural Actor, the first video-based method for learning a photorealistic neural human model that can be relighted, allows appearance editing, and can be controlled by arbitrary skeletal poses. Importantly, for learning our human avatar, we solely require a multi-view recording of the human under a known, but static lighting condition. To achieve this, we represent the geometry of the actor with a drivable density field that models pose-dependent clothing deformations and provides a mapping between 3D and UV space, where normal, visibility, and materials are encoded. To evaluate our approach in real-world scenarios, we collect a new dataset with four actors recorded under different light conditions, indoors and outdoors, providing the first benchmark of its kind for human relighting, and demonstrating state-of-the-art relighting results for novel human poses.
Figure 1: Our method learns a neural representation of an actor that is driven only by a 3D skeletal pose and allows rendering and editing of the actor’s appearance under new lighting conditions not seen during training. Our approach models pose-dependent clothing deformations, self-shadows, and provides a new way to map between 3D and UV spaces, enabling the intrinsic decomposition of dynamic sequences. At the same time, our method provides the possibility to easily edit the appearance and material properties at inference time. As input for training, our method only needs a few minutes of multi-view RGB videos and a light environment map.
@@ 1 Introduction
Creating a relightable and controllable virtual actor of a real person solely from RGB videos is an exciting and challenging topic that has numerous applications in virtual and augmented reality, gaming, human-machine interaction, and content creation. However, intrinsically decomposing the dynamic human geometry, global illumination, and material properties; achieving pose control with realistic dynamic deformations and photoreal relighting effects; and making the problem computationally tractable is still an unsolved problem in research.
@@
Relighting static scenes from multi-view imagery is a well-studied problem [49, 39, 17, 26]. However, directly applying such methods to moving humans, which are dynamic, highly articulated, and have complex material composition, is not trivially possible as these approaches do not account for dynamics and their training time ranges from multiple hours to even days, even for a single frame. Despite some works [14, 9, 29, 4] assuming a well-calibrated lightstage settings, only few human-specific works [6, 51, 13] try to achieve human intrinsic decomposition and relighting solely from multi-view video with static lights. However, at test time, they can solely replay the human motion under novel lighting conditions [6], or do not model global illumination and space-varying BRDF [13], which results in low visual fidelity.
@@
To address this, we propose the first approach that achieves intrinsic decomposition of dynamic human avatars, enabling pose-control and photorealistic rendering, while solely requiring a multi-view video of the real human under a single and static light condition during training (see Fig. 1). At inference time, we only need arbitrary driving skeletal pose and environment map. We solve this challenging problem by modeling the human with an implicit density field controlled via a human body model [27], which provides a pose-driven implicit geometry for arbitrary poses while maintaining a coarse correspondence between points in 3D space and the body model’s surface in the respective 2D UV map. In addition, we propose a novel intrinsic decomposition in 2D UV space rather than in 3D space, making the problem computationally tractable for the long and dynamic human sequences. Specifically, we represent normal and visibility information as dynamic maps, and albedo and roughness as static maps. Our key contribution is to sample normal and visibility from the posed geometry and aggregate them into sparse UV maps, which are then densified by the proposed NormalNet and VisibilityNet models. This allows our approach to model wrinkles and self-shadows even for long sequences. The albedo and roughness maps are optimized during training and sampled during rendering. To account for dynamic effects and spurious miss-alignments in the UV projection, this sampling process is guided by the proposed UVDeltaNet, which is intrinsically learned in our pipeline. Finally, given an environment map and a virtual camera position, our neural renderer produces the final image of the neural actor.
@@
To train and evaluate our method, we collected a new real dataset for novel pose and novel lighting human synthesis, composed of four actors recorded under six different light conditions by a multi-view camera setup, including indoor and outdoor environments and a light probe for each sequence, providing the first real in-the-wild benchmark for controllable human relighting. To summarize, our main technical contributions are as follows:
  • The first method to learn a relightable neural actor from multi-view RGB videos that allows for pose control, free-viewpoint rendering, diffuse and specular reflection decomposition, and material editing at inference time;
  • A new approach to compute dynamic normals and visibility by sparsely sampling a neural field and densifying the results with NormalNet and VisibilityNet, which makes the problem of pose-driven human relighting tractable;
  • UVDeltaNet, which accounts for the dynamic effects in UV space, allowing us to learn an explicit and static material representation that is easy to edit at inference time.
In our experimental evaluation, we compare our approach with two baseline variations and show the effectiveness of each component in our method with an ablation study. The quality of our results are extensively illustrated in the paper and in the supplementary materials.
@@ 2 Related Work
Relighting a controllable human model with implicit representations is still an open challenge in research. We summarize in Tab. 1 the most related methods, all of which are unable to perform our task under similar assumptions.
Method Diffuse Specular Pose Control Dynamic Deform. Eval. Real Data Editable Neural Actor [25] Relighting4D [6$\star$ RANA [13$\dagger$ Ours $\star$ Results do not show specular effects. $\dagger$ Source code not available.
Table 1: Compared to existing approaches, only our method can perform intrinsic decomposition with diffuse and specular effects, model dynamic deformations, and simultaneously supports pose control and editing. Our method is the first to evaluate human relighting numerically under new poses and new lights on real data.
@@ 2.1 Relighting Static Scenes
Classical approaches for relighting and editing perform intrinsic decomposition through inverse rendering [35]. With recent advances in neural rendering [40], light transportation can be taken into account with fully differentiable rendering pipelines. For instance, Zhang et al. [49] address the problem of recovering geometry and spatially-varying BRDF with an implicit scene representation. However, this requires solving the light transport by evaluating the NeRF model for each light ray, which is very inefficient. NeRV [39] and Zhang et al. [50] mitigate this problem with a neural visibility field that is jointly optimized with the NeRF model, and NeRD [5] adds further relaxation by ignoring cast shadows. However, this assumption cannot be made for humans, which are highly articulated and often experience prominent self-casting shadows. Jin et al. [17] proposes TensorIR, an efficient tensor-based representation that solves the inverse problem by casting camera- and secondary-rays. IRON [47] proposes a two-stage process for obtaining a mesh with material properties, and Lyu et al. [28] and NeRO [26] focus on multi-bounce illumination and reflective objects, which is usually not the case for humans. A common limitation of these methods is in the assumption of a static scene; therefore, they cannot be trivially adapted to dynamic and controllable humans.
@@ 2.2 Human Models with Pose Control
Modeling humans is a long-standing topic. Traditionally, meshes from an actor can be automatically rigged to a skeleton [7] using Linear Blending Skinning (LBS) [21]. Parametric human models offer pose and shape control [27, 33] and, more recently, implicit human bodies can be articulated [31, 3, 30]. Such models offer a generic and controllable representation but are unable to represent individual details. Template-based methods synthesize a realistic level of details [10, 11], but have to deal with non-linear pose-dependent deformations.
Several approaches model appearance and geometry with an implicit field anchored on a coarse human mesh. Neural Body [34] optimizes latent vectors anchored in the vertices of the human mesh and Neural Actor [25] incorporates texture maps as a conditioning to break down the mapping from pose to dynamic effects (one-to-many mapping). ARAH [44] models the surface of the human with an SDF, Dual-Space NeRF [52] learns the color and light in separate spaces, and Kwon et al. [19] propose a generalizable approach. Shysheya et al. [38] jointly learn textures and a neural renderer but fall short in reconstructing the 3D human body. A few works [46, 16] approach a restrictive scenario of reconstructing 3D humans from a monocular video but fail to model fine details under new poses. Common for these methods is the entanglement of material properties and global illumination, making them restricted to the same appearance as observed during training.
@@ 2.3 Relighting Human Models
Only a few methods in the literature tackle the problem of relighting humans from video data under static illumination. Li et al. [22] create a relightable human mesh by estimating a low-frequency environment using Spherical Harmonics (SH). Some approaches target the problem of single-image relighting [15] and static pose [51] under unknown lights, which is an extremely ill-posed problem due to the high ambiguity between albedo and lighting. Relighting4D [6] extends Neural Body [34] to estimate geometry and reflectance, but can only replay the same sequence observed during training under new lights and is not able to generate new poses. RANA [13] allows for pose control but assumes Lambertian reflectance, ignoring different material properties such as skin and clothes, and do not model cast shadows, which is a strong limitation for articulated humans.
In contrast, our method performs intrinsic decomposition with space-varying BRDF, models global illumination, allows for pose control and novel lights, and offers an easy manner to edit the material properties of the learned actor.
@@ 3 Method
Figure 2: Our method takes as input a 3D skeletal pose and a static environment map, and renders the neural actor from a virtual camera position. The pose-driven geometry (Sec. 3.2) learns an implicit density, from which we can compute normals and visibility. Our intrinsic decomposition disentangles normal, visibility, albedo and roughness maps in UV space (Sec. 3.3). The neural renderer (Sec. 3.4) outputs our prediction, which is supervised with the reference image. Yellow represents learnable components and green non-learnable components.
Our goal is to generate a digital human avatar (neural actor) from multi-view RGB images that can be relighted, pose-controlled and rendered from novel viewpoints at test time. We decompose the observed appearance into normals, visibility w.r.t. the light sources, and BRDF represented as albedo and roughness. With these components and an environment map, we employ the microfacet model [42] to render again the observed images. In addition—since we target a drivable actor—the geometry components need to be controllable by the skeletal pose.
Fig. 2 shows an overview of our method divided into three main parts. The first part is a pose-driven geometry model, where the goal is to obtain a skeleton-controllable implicit 3D density field that models the geometry of the actor with pose-dependent deformations. The second part is the new intrinsic decomposition model whose goal is to disentangle geometry represented as normal and visibility information, and material properties as albedo and roughness. Our key insight is to represent these components in UV space rather than in 3D coordinates and use the neural field as a guide in the mapping function between 3D and UV spaces. Finally, we integrate the microfacet model into our rendering pipeline that takes as input the disentangled components (i.e., normal, visibility, albedo, and roughness) and an environment map, and renders the image of the neural actor in an arbitrary pose from a free-viewpoint.
3.1 Inputs and Definitions
@@
During training, our method is supervised with multi-view calibrated video streams $\mathbf{V}_{c}=\{\mathbf{I}_{c,t}\}$, where ${c\in[1,\dots,C]}$ is the camera index and $\mathbf{I}_{c,t}$ is the image from camera $c$ at frame $t$. We obtain a foreground segmentation mask $\mathbf{M}_{c,t}$ for each video frame [23]. The input 3D skeletons drive the LBS model to obtain posed human meshes used to extract texture maps from the multi-view images [2]. These texture maps are used during training only. At inference time, they are predicted by a V2V [45] network in a similar way to Neural Actor [25].
3.2 Pose-driven Implicit Human Model
We represent human geometry as an implicit field guided by the skeletal pose parameters $\mathbf{P}_{t}$. We follow the architecture of Neural Actor [25] and optimize our model through volumetric rendering. However, differently from Liu et al. [25], we sample image patches during training, which allows our method to only process the triangles intersecting the patch region. From each pixel in a patch, we cast rays represented by ${\mathbf{r}_{n}=\mathbf{o}+\delta_{n}\mathbf{d}}$, with origin $\mathbf{o}\in\mathbb{R}^{3}$ in the optical center of the camera and direction $\mathbf{d}\in\mathbb{R}^{3}$. Each point in the ray is distant from the camera by the depth value $\delta_{n}$. The sampled points $\mathbf{r}_{n}$ are projected to the mesh surface, resulting in the local coordinates $(u_{n},v_{n},h_{n})$, where $(u,v)$ is the surface coordinate in the UV space and $h$ is the signed distance to the mesh. The corresponding points $\mathbf{x}_{n}=(x,y,z)$ in the canonical space are obtained by unwarping and normalizing the points in world coordinates following the inverse projection of the LSB model.
Next, DeformNet predicts a 3D deformation in the canonical space by taking as input the local coordinates and texture features $\psi_{n}$ extracted by a ResNet [12] backbone. This is important to model pose-dependent effects, such as wrinkles, that are not modeled by the LBS mesh. Finally, we produce an image
$\hat{\mathbf{I}}(\mathbf{r})=\sum_{n=1}^{N}w_{n}{c_{n}},\quad w_{n}=T_{n}\left(1-\text{exp}\left(-\sigma_{n}{\bar{\delta}_{n}}\right)\right),$ (1)
where $T_{n}=\text{exp}\left(-\sum_{j=1}^{n-1}\sigma_{j}{\bar{\delta}_{j}}\right)$, $w_{n}$ is the density weight from standard volumetric rendering [32], $\bar{\delta}_{j}=\delta_{j+1}-\delta_{j}$ is the distance between adjacent samples, and $c_{n}$ and $\sigma_{n}$ are the color and density predictions.
Since our goal at this stage is to obtain a drivable high-quality geometry, we only use the image $\hat{\mathbf{I}}$ for supervising the geometry model. The loss for this part is defined as:
$\mathcal{L}_{\text{geo}}=\lambda_{\text{L2}}\mathcal{L}_{\text{L2}}+\lambda_{\text{vgg}}\mathcal{L}_{\text{vgg}}+\lambda_{\text{hard}}\mathcal{L}_{\text{hard}}+\lambda_{\sigma}\mathcal{L}_{\sigma},$ (2)
where $\mathcal{L}_{\text{L2}}$ and $\mathcal{L}_{\text{vgg}}$ are the L2 and perceptual losses  [20] between $\mathbf{I}$ and $\hat{\mathbf{I}}$, $\mathcal{L}_{\text{hard}}$ is the hard loss from Rebain et al. [36], and $\mathcal{L}_{\sigma}$ pushes the density in the empty space to zero. Specifically, the density defined as $\sigma=\text{ReLU}(\sigma^{\prime})$ suffers from zero-gradient on empty regions and the pre-activated $\sigma^{\prime}$ is pushed towards negative values with
$\mathcal{L}_{\sigma}=\left(1-\mathbf{M}(\mathbf{r})\right)\text{sigmoid}(\sigma^{\prime}),$ (3)
where $\mathbf{M}(\mathbf{r})$ is the binary mask at the pixel intersected by $\mathbf{r}$. $\lambda_{(.)}$ are defined empirically and set to $\lambda_{\text{L2}}=100$ and $\lambda_{\text{vgg}}{=}\lambda_{\sigma}{=}\lambda_{\text{hard}}{=}0.01$ in all experiments. After training, we discard the color from Eq. 1 and only use the density $\sigma$ to implicitly represent the geometry of the neural actor.

### v2

# Project page: https://vcai.mpi-inf.mpg.de/projects/RNA

@@ Abstract
Creating a controllable and relightable digital avatar from multi-view video with fixed illumination is a very challenging problem since humans are highly articulated, creating pose-dependent appearance effects, and skin as well as clothing require space-varying BRDF modeling. Existing works on creating animatible avatars either do not focus on relighting at all, require controlled illumination setups, or try to recover a relightable avatar from very low cost setups, i.e. a single RGB video, at the cost of severely limited result quality, e.g. shadows not even being modeled. To address this, we propose Relightable Neural Actor, a new video-based method for learning a pose-driven neural human model that can be relighted, allows appearance editing, and models pose-dependent effects such as wrinkles and self-shadows. Importantly, for training, our method solely requires a multi-view recording of the human under a known, but static lighting condition. To tackle this challenging problem, we leverage an implicit geometry representation of the actor with a drivable density field that models pose-dependent deformations and derive a dynamic mapping between 3D and UV spaces, where normal, visibility, and materials are effectively encoded. To evaluate our approach in real-world scenarios, we collect a new dataset with four identities recorded under different light conditions, indoors and outdoors, providing the first benchmark of its kind for human relighting, and demonstrating state-of-the-art relighting results for novel human poses.
@@ 1 Introduction
Creating a relightable and controllable virtual actor of a real person solely from multi-view videos under static illumination is an exciting and challenging topic that has numerous applications in virtual and augmented reality, gaming, human-machine interaction, and content creation. However, it requires modeling the outgoing radiance of a particular surface point on the human, which depends on the pose-dependent and highly non-rigid surface deformation of the human body, the complex and spatially varying materials, as well as the scene lighting. Intrinsically decomposing these quantities solely from multi-view video and enabling novel compositions that are of highly photorealistic quality defines an unsolved and challenging inverse problem in the literature.
Figure 1: Our method learns a neural actor that is driven only by a 3D skeletal pose and allows rendering and editing of the actor’s appearance under new lightning and poses not seen during training. Our approach models pose-dependent deformations, self-shadows, and performs intrinsic decomposition of dynamic humans through a pose-dependent mapping between 3D and UV spaces, which also enables editing the appearance and material properties at inference time. For training, our method only needs a multi-view video and an environment map.
@@
Relighting static scenes from multi-view imagery is a well-studied problem [63, 49, 22, 33]. However, directly applying such methods to moving humans, which are dynamic, highly articulated, and have complex material composition, is not trivially possible, as these approaches do not account for dynamics. On the contrary, some works have focused on building drivable and dynamic full-body avatars from multi-view video [32, 42, 55] without considering relightability at all. Despite some works [18, 11, 36, 4, 67] assuming a well-calibrated lightstage settings, only few human-specific works [6, 65, 17] try to achieve human intrinsic decomposition and relighting from videos with static lights. While they are able to recover a relightable human representation from a single RGB video, they can solely replay the human motion under novel lighting conditions [6], or do not model self-shadows and space-varying BRDF [17], which results in low visual fidelity. In stark contrast, we assume a multi-view setting while accounting for high-frequency surface deformations as well as induced high-frequency lighting effects such as self-shadows and shading, which are indispensable for a truly immersive experience.
To address this, we propose Relightable Neural Actor, a new method that achieves intrinsic decomposition of dynamic human avatars, enabling pose-control, material editing, and realistic rendering with pose-dependent effects, while solely requiring a multi-view video of the real human under a single and static light condition during training (see Fig. 1). At inference time, we only need an arbitrary driving skeletal pose and an environment map.
@@
We solve this challenging problem by leveraging an implicit density field from a pose-driven neural actor and by proposing a novel intrinsic decomposition guided by the implicit density field. This strategy allows our method to model pose-dependent deformations in the geometry as dynamic maps, while representing the albedo and roughness properties of the actor as static maps. Our approach not only makes the problem computationally tractable for the long and dynamic human sequences, but also enables editing the neural actor’s appearance and materials at inference time. Our key contribution is to sample normal and visibility from the posed density field and aggregate them into sparse UV maps, which are then densified by the proposed NormalNet and VisibilityNet models. This allows our approach to effectively model wrinkles and self-shadows even for long sequences. The albedo and roughness maps are optimized during training and sampled by the proposed UVDeltaNet, which accounts for pose-dependent deformations in the UV map and is intrinsically learned in our pipeline. Finally, given an environment map and a virtual camera position, our neural renderer produces the final image of the neural actor.
@@
To train and evaluate our method, we collected a real-world dataset for novel pose and novel lighting human synthesis, composed of four actors recorded under six different light conditions by a multi-view camera setup, including indoor and outdoor environments and a light probe for each sequence, providing the first real in-the-wild benchmark for controllable human relighting.
To summarize, our main technical contributions are as follows: 1) The first method to learn a relightable neural actor from multi-view video under static lights that allows for pose control, appearance editing, and free-viewpoint rendering under arbitrary illumination while modeling pose-dependent effects. 2) We propose an intrinsic decomposition of diffuse and specular reflection based on a new approach to compute pose-dependent normals and visibility by sparsely sampling a neural field and densifying the results with NormalNet and VisibilityNet, making the problem of pose-driven human relighting tractable. Our approach relies on a mapping between 3D and UV space implicitly learned by UVDeltaNet, allowing us to learn an explicit and static material representation that is easy to edit at inference time.
In our experimental evaluation, we compare our approach with two baselines in the proposed real data benchmark, showing the superiority of our approach in synthesizing relightable humans on challenging motion sequences. The effectiveness of our method is demonstrated by ablation studies and qualitative results.
@@ 2 Related Work
Since the goal of this work is to build a full-body relightable and animatible avatar, we will not discuss works [44, 57, 28, 46, 9, 50, 10, 60, 59, 47] focusing on face relighting as both settings have different requirements and challenges.
@@ 2.1 Relighting Static Scenes
Classic approaches for relighting and editing perform intrinsic decomposition through inverse rendering [43]. With recent advances in neural rendering [51], light transportation can be taken into account with fully differentiable rendering pipelines. For instance, Zhang et al. [63] address the problem of recovering geometry and spatially-varying BRDF with an implicit scene representation. However, this requires solving the light transport by evaluating the NeRF model for each light ray, which is very inefficient. NeRV [49] and Zhang et al. [64] mitigate this problem with a neural visibility field that is jointly optimized with the NeRF model, and NeRD [5] adds further relaxation by ignoring cast shadows. However, this assumption cannot be made for humans, which are highly articulated and often experience prominent self-casting shadows. Jin et al. [22] propose TensorIR, an efficient tensor-based representation that solves the inverse problem by casting camera and secondary rays. IRON [61] proposes a two-stage process for obtaining a mesh with material properties, and Lyu et al. [35] and NeRO [33] focus on multi-bounce illumination and reflective objects, which is usually not the case for humans. The methods mentioned above share the assumption of a static scene and cannot be trivially adapted to dynamic and controllable humans.
@@ 2.2 Human Models with Pose Control
Modeling humans is a long-standing topic. Traditionally, meshes from an actor can be automatically rigged to a skeleton [7] using Linear Blending Skinning (LBS) [27]. Parametric human models offer pose and shape control [34, 41] and, more recently, implicit human bodies can be articulated [37, 3, 38]. Such models offer a generic and controllable representation, but are unable to represent individual details. Template-based methods synthesize a realistic level of details [14, 15, 21], but have to deal with non-linear pose-dependent deformations.
Several approaches model appearance and geometry with an implicit field anchored on a coarse human mesh. Neural Body [42] optimizes latent vectors anchored in the vertices of the human mesh and Neural Actor [32] incorporates texture maps as a conditioning to break down the mapping from pose to dynamic effects (one-to-many mapping). ARAH [55] models the human surface with an SDF, Dual-Space NeRF [66] learns the color and light in separate spaces, and Kwon et al. [24] propose an approach that generalizes to new identities and poses. Shyshey et al. [48] jointly learn textures and a neural renderer but fall short in reconstructing the 3D human body. A few works [58, 20] approach a restrictive scenario of reconstructing 3D humans from a monocular video but fail to model fine details under new poses. Common to all these type of methods [13, 12, 25, 40] is the entanglement of material properties and illumination, making them restricted to the same appearance as observed during training.
@@ 2.3 Relighting Human Models
Only a few methods in the literature tackle the problem of relighting humans from videos under static illumination. Li et al. [29] create a relightable human mesh by estimating a low-frequency environment using Spherical Harmonics. Some approaches target the problem of single-image [19] and static pose [65] relighting under unknown lights, which is an extremely ill-posed problem due to the high ambiguity between albedo and lighting. Relighting4D [6] extends Neural Body [42] to estimate geometry and reflectance, but can only replay the same sequence observed during training under new lights and is not able to generate new poses. RANA [17] allows for pose control but assumes Lambertian reflectance, ignoring different material properties such as skin and clothes, and does not model cast shadows, resulting in unrealistic renderings for articulated humans. In contrast, our method performs intrinsic decomposition with space-varying BRDF, models illumination and self-shadows, allows for pose control and novel lights, and offers an easy manner to edit the material properties of the learned neural actor.
@@ 3 Method
Our goal is to learn a digital human avatar (neural actor) from RGB images that models pose-dependent effects and can be relighted, pose-controlled, and rendered from novel viewpoints at test time. To achieve relighting, we decompose the observed appearance into normals, visibility with respect to each light source, and BRDF represented as albedo and roughness. With these components and an environment map, we employ the microfacet model [53] to render the output image of our method, which is compared to the observed images. In addition—since we target a drivable actor—the geometry components need to be controllable by the skeletal pose.
Fig. 2 shows an overview of our method divided into three main components. The first is a pose-driven implicit geometry model, where the goal is to obtain a skeleton-controllable implicit 3D density field that models the geometry of the actor with pose-dependent deformations. This model is adapted from Neural Actor [32] and is depicted here since our approach builds upon its internal design. Note that Neural Actor does not perform intrinsic decomposition and is unable to perform relighting or appearance editing, which is our main goal. The second component is our proposed intrinsic decomposition approach whose goal is to disentangle geometry represented as normal and visibility information, and material properties as albedo and roughness. Our key insight is to represent these components in UV space rather than in 3D coordinates, using the density field as guidance in the mapping function between 3D and UV spaces. This allows our method to model complex deformations that are pose-dependent in a compact UV map that can be easily computed and queried. Finally, we integrate the microfacet model [53] into our rendering pipeline that takes as input the disentangled components (i.e., normal, visibility, albedo, and roughness) and an environment map, and renders the image of the neural actor in an arbitrary pose and light condition from a free-viewpoint.
Figure 2: Our method takes as input a 3D skeletal pose and a static environment map, and renders the neural actor from a virtual camera position. The pose-driven geometry (Sec. 3.1) learns an implicit density function, from which we compute the normal and visibility information. Our intrinsic decomposition disentangles normal, visibility, albedo and roughness maps in UV space (Sec. 3.2). The neural renderer (Sec. 3.3) outputs our prediction, which is supervised with the reference image. Yellow represents learnable components and green non-learnable components.
3.1 Preliminaries
3.1.1 Human Pose and Environment Map.
@@ 3.1 Preliminaries
3.1.2 Training Data.
During training, our method is supervised with multi-view calibrated video streams $\mathbf{V}_{c}=\{\mathbf{I}_{c,t}\}$, where ${c\in\{1,\dots,C\}}$ is the camera index and $\mathbf{I}_{c,t}$ is the image from camera $c$ at frame $t$. We obtain a foreground mask $\mathbf{M}_{c,t}$ for each video frame [30]. The coarse human mesh and 3D skeletons can be obtained with multi-view pose tracking [1]. The input 3D skeletons drive the LBS model to obtain posed human meshes used to extract texture maps from the multi-view images [2]. These texture maps are used during training to learn pose-dependent deformations on the implicit human model. At inference time, a V2V [56] network predicts texture maps from posed normal maps [32].
3.1.3 Implicit Human Model.
For realistically relighting articulated avatars, we need to model human geometry with deformations caused by dynamic and varied poses. Therefore, we follow Neural Actor [32] to represent the pose-driven geometry as an implicit field guided by the 3D skeletal pose parameters and conditioned by texture maps, which provide the ability to model high-frequency deformations. Specifically, a standard NeRF model [39] is optimized through volumetric rendering in the canonical space of the coarse human mesh and the canonical space is deformed based on ResNet [16] features extracted from the texture maps, which provide the ability to represent detailed wrinkles and deformations not modeled by the coarse mesh. During optimization, from each pixel in the reference images we cast a ray ${\mathbf{r}_{n}=\mathbf{o}+\delta_{n}\mathbf{d}}$, where $n={1,\dots N}$ is the index of each point in the ray, $\mathbf{o}\in\mathbb{R}^{3}$ is the origin defined in the optical center of the camera, $\mathbf{d}\in\mathbb{R}^{3}$ is the ray direction, and $\delta_{n}$ is the depth (distance to the camera). The sampled points $\mathbf{r}_{n}$ are projected to the mesh surface, resulting in the local coordinates $(u_{n},v_{n},h_{n})$, where $(u,v)$ is the surface coordinate in the UV space and $h$ is the signed distance to the mesh.